# Lorentz-violating extension of Wigner function formalism and chiral kinetic theory

### v1

@@ VI Discussions
We studied the fermionic part of the extended QED which is the modified Dirac equation with a set of LSV terms, in the presence of external electromagnetic fields. It is observer Lorentz symmetric as a reminiscent of the spontaneously broken Lorentz invariance of the fundamental theory which may be String Theory. We proceed by quantizing spinor fields using standard methods and utilizing these spinor operators to construct a relativistic Wigner function through conventional quantum field theory techniques.
The quantum kinetic equation governing the Wigner function is accomplished by following the well-established formulation of [2]. By decomposing the Wigner function in terms of the Clifford algebra generators constructed from the gamma matrices, we derive a set of coupled equations satisfied by the scalar, pseudoscalar, vector, axial-vector and antisymmetric tensor fields. Notably, in the the massless limit, the equations of the vector and axial-vector fields are decoupled from the rest. In fact, we focus on the chiral fermions and adopt the semiclassical approximation where the terms up to the first order in Planck constant are considered.
@@
In principle, one could derive semiclassical kinetic equations for massive fermions by examining the defining equations (9)-(13). However, even in the case of ordinary Dirac fermions, this task is intricate, as evidenced by the complexities encountered when employing the approaches outlined in [25] or [26].
Integration of (78) over the zeroth component of momentum to obtain the nonrelativistic (3D) chiral kinetic theory is a desired step, albeit complicated by the presence of $G_{\mu\nu}$. A pragmatic solution involves setting $a_{0}=0,\ G_{0\mu}=\delta_{0\mu}$, yielding $\bar{p}_{0}=p_{0},$ which facilitates to get the mass-shell condition dictated by the Dirac delta function.
By integrating the equations governing the components of the Wigner function, (9)-(13), with respect to the zeroth component of momentum, one can derive what is known as the equal-time formulation, a technique elucidated in detail in [27, 28, 29, 30].

### v2

@@ VI Discussions
We studied the fermionic part of the extended QED which is the modified Dirac equation with a set of LSV terms, in the presence of external electromagnetic fields. It is observer Lorentz symmetric as is reminiscent of the spontaneously broken Lorentz invariance of the fundamental theory which may be string theory. We proceed by quantizing spinor fields using standard methods and utilizing these spinor operators to construct a relativistic Wigner function through conventional quantum field theory techniques.
The quantum kinetic equation governing the Wigner function is accomplished by following the well-established formulation of [2]. By decomposing the Wigner function in terms of the Clifford algebra generators constructed from the $\gamma$ matrices, we derive a set of coupled equations satisfied by the scalar, pseudoscalar, vector, axial-vector and antisymmetric tensor fields. Notably, in the massless limit, the equations of the vector and axial-vector fields are decoupled from the rest. In fact, we focus on the chiral fermions and adopt the semiclassical approximation where the terms up to the first order in Planck constant are considered.
@@
How can one experimentally test the results of the modified chiral kinetic theory? This may be achieved through the currents (71) and (72), which yield the CME and CSE. In high-energy physics, experimental evidence for the CME is investigated in heavy-ion collisions. The current experimental status has been recently reviewed in [27]. We have seen that LSV parameters modify the coefficient of the magnetic field (75), by $|G|^{-1},$ which would be nearly impossible to detect. However, note that the modified current is along $\bar{\bm{B}},$ hence the direction of the CME is altered by LSV parameters. Although this effect is small, it is the unique source in the deviation in the direction. This can even help in experimentally observing the CME, which has been investigated only along the direction of the magnetic field.
In principle, one could derive semiclassical kinetic equations for massive fermions by examining the defining equations (9)-(13). However, even in the case of ordinary Dirac fermions, this task is intricate, as evidenced by the complexities encountered when employing the approaches outlined in [28, 29].
Integration of (86) over the zeroth component of momentum to obtain the nonrelativistic (3D) chiral kinetic theory is a desired step, albeit complicated by the presence of $G_{\mu\nu}$. A pragmatic solution involves setting $a_{0}=0,\ G_{0\mu}=\delta_{0\mu}$, yielding $\bar{p}_{0}=p_{0},$ which facilitates to get the mass-shell condition dictated by the Dirac $\delta$ function.
By integrating the equations governing the components of the Wigner function (9)-(13), with respect to the zeroth component of momentum, one can derive what is known as the equal-time formulation, a technique elucidated in detail in [30, 31, 32, 33].